x^2-6x=16
Solve by completing the square.
this is what I did...but it isn't the correct way to do it.
(x-8)(x+2)
x=8, x=-2

Answer by nerdybill(7384)

This site explains the entire process:
http://www.purplemath.com/modules/sqrquad.htm
.
x^2-6x=16
Divide the -6 by 2 and square it to get 9. Add it to both sides:
x^2-6x+9=16+9
(x-3)^2 = 25
take the square root of both sides:
x-3 = (+-)5
x = 3(+-)5
.
x = {-2, 8}